40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 16:46:31 2019, gautam, Update, IOO, IMC not locking PCdrive_RMS.png
    Reply  Tue Jun 4 10:56:31 2019, gautam, Update, IOO, IMC diagnostics IMCdiag.pdf
       Reply  Thu Jun 6 22:11:53 2019, Koji, Update, IOO, IMC diagnostics 
          Reply  Thu Jun 13 18:01:18 2019, aaron, Update, IOO, IMC diagnostics 
             Reply  Fri Jun 14 13:10:00 2019, aaron, Update, IOO, IMC diagnostics 
                Reply  Mon Jun 17 12:37:16 2019, aaron, Update, IOO, IMC diagnostics 
                   Reply  Tue Jun 18 20:35:07 2019, aaron, Update, IOO, IMC diagnostics 
                      Reply  Wed Jun 19 19:12:51 2019, aaron, Update, IOO, IMC diagnostics PMC_Error_Spectrum.pdfPDH_signal.pdfPDH_signal_full.pdf
                         Reply  Mon Jun 24 15:49:05 2019, aaron, Update, IOO, IMC diagnostics sweep_config_updated.png
                            Reply  Tue Jul 2 19:36:18 2019, aaron, Update, IOO, IMC diagnostics PDH_error.pdfPMC_Error_Spectrum.pdf
          Reply  Fri Jun 21 19:36:26 2019, Koji, Update, IOO, IMC diagnostics IMC_error.pdf
             Reply  Sun Jun 23 08:09:53 2019, gautam, Update, IOO, NPRO diagnostics consolidatedOscopeScreenCaps.pdfRINcomp.pdfPLL_OLTF.pdfPLLnoise.pdf
                Reply  Sun Jun 23 09:36:32 2019, gautam, Update, IOO, IMC is locking normally again PMCdiag.pdf
                   Reply  Sun Jun 23 14:43:14 2019, Koji, Update, IOO, IMC is locking normally again 
                   Reply  Mon Jun 24 08:12:10 2019, gautam, Update, IOO, IMC is locking normally again IMCdutycycle.png
                      Reply  Mon Jun 24 11:48:35 2019, gautam, Update, IOO, IMC in-loop error spectra and OLTF IMCfreqNoise.pdfIMC_OLTF.pdfIMC_CMboard.jpg
Message ID: 14686     Entry time: Fri Jun 21 19:36:26 2019     In reply to: 14659     Reply to this: 14687
Author: Koji 
Type: Update 
Category: IOO 
Subject: IMC diagnostics 

The IMC REFL error signal was measured to compare it with the other spectra (if we have).

The blue curve is the in-loop IMC error and the red is the dark noise. So they are not an apple-to-apple comparison. But the red noise is going to be suppressed by the loop, and still the red is below blue. This means that the blue curve is the measured noise rather than the readout noise.

We suspect that the current issue is the PC drive saturation (as usual). Does this indicate that the laser freq noise is actually increased?

----

Another suspect was that the degradation of the LO level. We used to have the issue of slowly dying ERA-5 (ERA-5SM indeed). The RF levels on the demod board were measured using an active probe.

The LO input: 0dBm, ERA-5 input: -2.7dBm and -2.1dBm for I and Q. I found that the outputs of the ERA-5SM were +10.5dBm and +10.6dBm.
This lead me to replace the chips but the situation was not changed. Then I realized that the LO levels should have been measured with the load replaced from the mixers to a 50Ohm load. Somehow these mixers lower the apparent LO levels. So I decided to say this is OK.

Attachment 1: IMC_error.pdf  1.030 MB  Uploaded Fri Jun 21 20:42:39 2019  | Hide | Hide all
IMC_error.pdf
ELOG V3.1.3-